perla is now 5 times as old as Marilou. in 10 years, perla is will be 3 times as old as marilou. how old is each.
Answer by mouk(232)

Let x = age of Marilou now
Then "perla is now 5 times as old as Marilou"
so now, Perla's age is 5x

In 10 years time, Marilou age = x+10
In 10 years time, Perla's age = 5x+10
"in 10 years, perla is will be 3 times as old as marilou"
so 5x+10 = 3(x+10)
5x+10 = 3x+30
2x = 20
x = 10
"how old is each"
Marilou's age is x so Marilou is 10 years old
Perla's age is 5x so Perla is 50 years old.
Check:
In 10 Years time:
Marilou is 20 years old
Perla is 60 years old
And 20 x 3 = 60